OpenStudy (anonymous):

if 15 horses are racing, and you randomly select 3 to be in your trifecta (1st,2nd,3rd pick to win), what is the probability that you will win?

OpenStudy (anonymous):

P(A AND B AND C) = (1/15)*(1/14)*(1/13) @satellite73 can you check if this is correct?

OpenStudy (kropot72):

The probability of the pick for first coming in first is 1/15. If the pick for first comes in first the probability of the pick for second coming in is 1/14. If the picks for first and second come in the probability of the pick for third coming in third is 1/13. The three events are independent. Therefore the probability of winning the trifecta is \[\frac{1}{15}\times \frac{1}{14}\times \frac{1}{13}\]